OpenStudy (anonymous):

Use the quadratic formula to find the zeros of the function. Round to tenths if necessary. y = 3x^ + 5x A. {-2.23, 1.7} B. {-1.7, 0} C. {-2.8, 1.7} D. no real solution

OpenStudy (ash2326):

@PsYcHoEmOmAtH Do you know the quadratic formula?

OpenStudy (campbell_st):

this is the formula for a quadratic f(x) = ax^2 + bx +c \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] in your question a = 3 and b = 5 and c = 0 substitute the values then evaluate to find the zeros

OpenStudy (anonymous):

I know that I got lost though*

OpenStudy (campbell_st):

4ac will dissapear as 4 x 3 x 0 = 0 you will be left with \[x = \frac {-5\pm \sqrt{5^2}}{6}\] so the 2 solutions are \[x=\frac{-5 - \sqrt{25}}{6}\] and \[x=\frac{-5+\sqrt{25}}{6}\] just evaluate these for the zeros